# How Global Calibration Strengthens Multiaccuracy

### v1

@@ 4.2 Strong Agnostic Learning from Calibrated Multiaccuracy
In this section, we show that provided the predictor $p$ that is $(\mathcal{C},\tau)$-multiaccurate is also well-calibrated, then $\mathrm{sign}(2p-1)$ in fact is a strong agnostic learner. Interestingly, our proof decomposes the roles of multi-accuracy and calibration neatly and highlights the contribution that both these properties make in order to obtain a strong agnostic learning guarantee. We will assume in this section that the hypothesis class $C$ contains functions from $\mathcal{X}\to[-1,1]$; as discussed in Section 3.2, this is essentially without loss of any generality.
@@
$\geq\operatorname*{\mathbf{E}}_{\mathcal{D}}[|2p(\mathbf{x})-1|]-2\operatorname*{\mathbf{E}}[|\operatorname*{\mathbf{E}}[\mathbf{y}|p(\mathbf{x})]-p(\mathbf{x})|]$
@@
The proof of Theorem 4.5 below follows immediately from Lemmas 4.3 (which applies to $\tau$-calibrated predictors) and 4.4 (which applies to $(\mathcal{C},\tau)$-multiaccurate predictors). This shows provided that for any $\tau>0$ it is possible to construct a predictor $p$ that is $(\mathcal{C},\tau)$-multiaccurate for $\mathcal{D}$ such that $p$ is also $\tau$-calibrated, then $\mathrm{sign}(2p-1)$ yields a strong agnostic learning guarantee for the class $\mathcal{C}$ under $\mathcal{D}$.
@@ 5 Hardcore Measures with Optimal Density
In this section we consider a domain $\mathcal{X}$ endowed with a base distribution $\mathcal{D}_{\mathcal{X}}$. We encourage the reader to think of $\mathcal{X}=\{0,1\}^{n}$ and $\mathcal{D}_{\mathcal{X}}$ to be the uniform distribution. A measure on $\mathcal{X}$ is a function $\mu:\mathcal{X}\rightarrow[0,1]$, which is not identically 00. A measure $\mu$ defines a probability distribution $\bar{\mu}$ by setting
@@
$\Pr_{x\sim\mathcal{D}_{\mathcal{X}}}[c(\mathbf{x})=g(\mathbf{x})]\leq 1-\delta.$
@@
Definition 5.2 (Relative complexity of a function and a function class [28, Definition 6]).
@@ 5.1 Weighted Multiaccuracy and Hardcore Measures
$\min(\eta,1-\eta)+\mathrm{cor}_{\bar{\mu}}(g,\mathcal{C})\leq\max_{c\in\mathcal{C}}\Pr_{\mathbf{x}\sim\mathcal{D}_{\mathcal{X}}}[c(\mathbf{x})=g(\mathbf{x})]\leq\max(\eta,1-\eta)+\mathrm{cor}_{\bar{\mu}}(g,\mathcal{C}).$ (9)
@@
Recall that $\bar{c}=1-c$ and that we are assuming that $\mathcal{C}$ is closed under complemntation. So we have
@@
To prove the lower bound, in order predict $g$ with accuracy $\min(\eta,1-\eta)+|\mathrm{cor}_{\bar{\mu}}(g,c)|$, we use $c$ if $\mathrm{cor}_{\bar{\mu}}(g,c)\geq 0$; else we use $\bar{c}$, which also lies in $\mathcal{C}$ by assumption. Maximizing over $c\in\mathcal{C}$ gives the claimed bound. ∎
@@
Lemma 5.7 can be seen as a generalization of Yao’s Lemma relating pseudorandomness and unpredictability (see Lemma 2.14 in [8] for a formal statement).
@@
$=\operatorname*{\mathbf{E}}_{x\sim\mathcal{D}_{\mathcal{X}}}[c(\mathbf{x})w(p(\mathbf{x}))\cdot|g(\mathbf{x})-p(\mathbf{x})|\cdot(2g(\mathbf{x})-1)]$
@@
$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{D}_{\mathcal{X}}}[w(p(\mathbf{x}))(g(\mathbf{x})-p(\mathbf{x}))]|=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\bar{\mu}_{w}}[(2g(\mathbf{x})-1)]\cdot\mathsf{dns}(\mu)=(2\eta-1)\cdot\mathsf{dns}(\mu).$
@@
$\left|\eta-\frac{1}{2}\right|\leq\frac{\varepsilon}{2\mathsf{dns}(\mu)}.$
@@
$\mathrm{cor}_{\bar{\mu}_{w}}(\mathcal{C},g)\leq\frac{\varepsilon}{\mathsf{dns}(\mu)}.$
@@
$\leq\frac{1}{2}+\frac{\epsilon}{2\mathsf{dns}(\mu)}+\mathrm{cor}_{\bar{\mu}_{w}}(\mathcal{C},g)$
$\leq\frac{1}{2}+\frac{3\varepsilon}{2\mathsf{dns}(\mu)}$
@@ 5.2 Optimal Density Analysis Using Calibration
We will use the following identity for $v\in[0,1]$ and $y\in\{0,1\}$ which can be viewed as the multilinear expansion of $|y-v|$:
$|y-v|=y(1-2v)+v.$
Using this with $y=g(x)$ and $t=p(x)$ we can write
@@
Since $w(p)\in[1,2]$ and $1-2p\in[-1,1]$, we have $|w(p)(1-2p)|\leq 2$. Since $p$ is $\tau$-calibrated, by Equation (4) applied to $v(p)=w(p)(1-2p)/2$, the predictor $p$ satisfies the bound
@@
$\leq\operatorname*{\mathbf{E}}_{\mathcal{D}^{g}}[(p(\mathbf{x})\cdot\mathbf{1}[p(\mathbf{x})<1/2]+(1-p(\mathbf{x}))\cdot\mathbf{1}[p(\mathbf{x})\geq 1/2])]+\operatorname*{\mathbf{E}}_{\mathcal{D}^{g}}[|p(\mathbf{x})-\operatorname*{\mathbf{E}}[\mathbf{y}|p(\mathbf{x})]|]$
@@
Often, the assumption of bounded $\mathsf{ECE}$ can be replaced with the weaker assumption that $\operatorname*{\mathbf{E}}[v(p)(y-p)]$ is bounded for specific functions $v$. In the proof of Lemma 5.10, we need to bound the calibration error for $v(p)=w(p)(1-2p)/2$, while in the proof of Lemma 5.11, we only needed to bound the error for $v(p)=2\cdot\mathbf{1}[p\geq 1/2]-1$.
@@ 5.3 Hardcore Measures with Optimal Density
Next we consider the algorithmic complexity of obtaining a predictor that satisfies both calibration and $(w_{\mathsf{Max}},\mathcal{C})$-weighted multiaccuracy, assuming an oracle for $(\alpha,\beta)$ weak agnostic learning for $\mathcal{C}$. Following previous work, we will consider the number of invocations to the weak agnostic learner. The key point is that since the only care about a single weight function $w_{\mathsf{Max}}$ which is bounded by $2$, the complexity of the weighted versions of multiaccuracy and calibrated multiaccuracy are the same as their unweighted counterparts. The work of [16] which introduced calibrated multiaccuracy shows that its complexity is comparable to that of multiaccuracy, and much lower than multicalibration [16, Section 7]. This relies on ideas that are by now standard in the multicalibration literature [24, 19, 16], so we will provide a sketch.
@@
To achieve calibration in addition to weighted multiaccuracy, we use the analysis of the calibrated multiaccuracy algorithm from [16, Section 7]. The key observation there is that if the calibration error is large, then recalibrating the predictor reduces the squared loss. Hence we can interleave recalibration steps with the calls to the weak learner, all the while ensuring that the squared loss decreases, and the number of calls to weak learner stays bounded by $O(1/\beta^{2})$.
Lastly, we use Theorem 5.13 to derive IHCL with optimal hardcore density $2\delta$. We use the calibrated multiaccuracy algorithm from [16, Section 7] to bound the complexity $\mathcal{C}_{t,q}$ of the class on which the input function $g$ is assumed to be $\delta$-hard.
@@
Let $\mathcal{C}$ be a family of functions $c:\mathcal{X}\rightarrow\{0,1\}$, let $\mathcal{D}_{\mathcal{X}}$ be a probability distribution over $\mathcal{X}$, and let $\epsilon,\delta,\tau>0$. There exist $t=O((1/(\epsilon^{2}\delta^{2})+1/\tau)\cdot\log(|\mathcal{X}|/\epsilon))$, $q=O((1/(\epsilon^{2}\delta^{2}))$ such that the following holds: If $g:\mathcal{X}\rightarrow\{0,1\}$ is $(\mathcal{C}_{t,q},\delta)$-hard on $\mathcal{D}_{\mathcal{X}}$, then there is a measure $\mu$ satisfying:
@@
We call the $\mathsf{calMA}$ algorithm from [16] with $\mathcal{C},\mathcal{D}^{g}$, weight function $w_{\mathsf{Max}}$, and desired multiaccuracy error $\epsilon\delta$, and obtain a $(w_{\mathsf{Max}},\mathcal{C},\epsilon\delta)$-multiaccurate predictor $p$ for $\mathcal{D}^{g}$. Using this predictor $p$, the input function $g$, and the weight function $w_{\mathsf{Max}}$, we construct the measure $\mu_{\mathsf{Max}}$ as per Definition 5.4; namely, $\mu_{\mathsf{Max}}(x)=w_{\mathsf{Max}}(p(x))\cdot|g(x)-p(x)|$. By Theorem 5.13, this measure $\mu_{\mathsf{Max}}$ achieves the claimed hardness and optimal density stated in Theorem 5.14. For the $\mathsf{calMA}$ algorithm, we assume access to a proper $(\epsilon\delta,\epsilon\delta)$-weak agnostic learner for $\mathcal{C}$ under $\mathcal{D}_{\mathcal{X}}$; this satisfy the parameter requirements in [16, Alg. 2]. [^12] In [16], the total number of calls to an $(\alpha,\beta)$-weak agnostic learner is analyzed. Here we are not concerned with the efficiency of the learner so we can assume a strong learner giving $\alpha=\beta$.
By Theorem 7.7 in [16], the total number of calls to the learner during the execution of the $\mathsf{calMA}$ algorithm is bounded by $O(1/(\epsilon\delta)^{2})$, which yields the claimed $q$ parameter. As for the total number of gates in the circuit, each iteration of the multiaccuracy step in the $\mathsf{calMA}$ algorithm requires a scalar multiplication, a finite-precision addition, and a projection onto the interval $[0,1]$. As argued in [8, Appx. A], this can be handled by $O(\log(1/\epsilon)+\log(1/|\mathcal{X}|)$ gates, where we perform computations up to a fixed precision of $\Theta(\epsilon)$ and $\log(1/|\mathcal{X}|)$ corresponds to the bit-length of the elements in $\mathcal{X}$. Lastly, the calibration-based steps in the $\mathsf{calMA}$ algorithm do not require further calls to the learner, but they require computing all of the values of the predictor $p$ ([16, Lemma 7.4]). We discretize $p$ with parameter $\tau$, so that the interval $[0,1]$ is partitioned into $O(1/\tau)$ intervals. To compute the $O(1/\tau)$ values of the predictor $p$, we can use a lookup table of size $O(1/\tau)$; such a lookup function can be computed by a circuit of size $O(1/\tau)$ [1, §4]. ∎
@@
[8] uses the multicalibration algorithm to obtain a stronger and more general version of Impagliazzo’s hardcore lemma, IHCL+++++ +. From it, they recover the IHCL theorem with density $2\delta$. However, because this recovery is done through the multicalibration algorithm, the $(t,q)$ parameters for the class $\mathcal{C}_{t,q}$, although polynomial in $1/\epsilon,1/\delta$, are much larger than the ones we have obtained in Theorem 5.14. In particular, using the multicalibration algorithm presented in [17], we require $q=O(1/(\epsilon^{6}\delta^{6}))$ calls to the weak agnostic learner [16]. When we recover the IHCL theorem from IHCL+++++ +, the parameters further increase to $q=1/(\epsilon^{2}\delta)^{6}$ oracle gates. [^13] This is because we can only glue together the pieces in the partition that are large enough and balanced enough; see [8, §3.3] for details.
@@
In particular, Impagliazzo’s boosting-based proof requires $q=O(1/(\epsilon^{2}\delta^{2}))$ oracle calls [27, Thm. 1], and all subsequent hardcore constructions retain the dependence on $1/\epsilon^{2}$ in $q$ [35, 26], including the uniform versions of the hardcore lemma [45, 5, 48].
The size loss of $O(1/\epsilon^{2})$ between the class $\mathcal{C}_{t,q}$ (on which we assume the input function $g$ is mildly hard) and the class $\mathcal{C}$ (on which $g$ is strongly hard) is inevitable in all boosting-based proofs of IHCL [37], Blanc, Hayderi, Koch, and Tan recently showed that this circuit size loss is necessary and that these parameters are in fact optimal and inherent to the hardcore lemma, regardless of proof strategy [6]. The dependence on $\delta$ can be improved: Nisan’s min-max proof in Impagliazzo’s original paper yields $q=O(1/\epsilon^{2}\log(1/(\epsilon\delta))$ [27, Lemma 2], and it was further improved to $q=O(1/\epsilon^{2}\log(1/\delta))$ by Kilvans and Servedio who showed an elegant connection to boosting [35, Thm. 25]. But these proofs do not give the optimal density.
Deriving IHCL from IHCL+++++ +.
The IHCL+++++ + theorem of [8] partitions the space $\mathcal{X}$ into level sets $S_{v}$. Within each level set, they construct a hardcore $\mathcal{H}_{v}$ distribution of density $2\min(v,1-v)$. We also know that $\operatorname*{\mathbf{E}}[\min(v,1-v)]\geq\delta$. To derive the IHCL with density $2\delta$, we need to glue these distributions together to get a measure of density $2\delta$. In their paper, [8] suggest the following gluing: sample $S_{v}$ according to $\mathcal{D}_{\mathcal{X}}$ and then sample within $S_{v}$ from $\mathcal{H}_{v}$. This gluing as stated might not give a distribution with bounded density. The reason being that when we take convex combinations of distributions over disjoint regions, the min-entropy of the combination is the maximum of the min-entropies, not the average. A simple fix is to go from hardcore distributions to sets, and then take the union of sets.

### v2

@@ 4.2 Strong Agnostic Learning from Calibrated Multiaccuracy
In this section, we show that provided the predictor $p$ that is $(\mathcal{C},\tau)$-multiaccurate is also well-calibrated, then $\mathrm{sign}(2p-1)$ in fact is a strong agnostic learner. Interestingly, our proof decomposes the roles of multiaccuracy and calibration neatly and highlights the contribution that both these properties make in order to obtain a strong agnostic learning guarantee. We will assume in this section that the hypothesis class $C$ contains functions from $\mathcal{X}\to[-1,1]$; as discussed in Section 3.2, this is essentially without loss of any generality.
@@
[…]
@@
The proof of Theorem 4.7 below follows immediately from Lemmas 4.5 (which applies to $\tau$-calibrated predictors) and 4.6 (which applies to $(\mathcal{C},\tau)$-multiaccurate predictors). This shows provided that for any $\tau>0$ it is possible to construct a predictor $p$ that is $(\mathcal{C},\tau)$-multiaccurate for $\mathcal{D}$ such that $p$ is also $\tau$-calibrated, then $\mathrm{sign}(2p-1)$ yields a strong agnostic learning guarantee for the class $\mathcal{C}$ under $\mathcal{D}$.
@@ 5 Hardcore Measures with Optimal Density
In this section we consider a domain $\mathcal{X}$ endowed with a base distribution $\mathcal{D}_{\mathcal{X}}$. We encourage the reader to think of $\mathcal{X}=\{0,1\}^{n}$ and $\mathcal{D}_{\mathcal{X}}$ to be the uniform distribution. A measure on $\mathcal{X}$ is a function $\mu:\mathcal{X}\rightarrow[0,1]$, which is not identically 0. A measure $\mu$ defines a probability distribution $\bar{\mu}$ by setting
@@
$\Pr_{\mathbf{x}\sim\mathcal{D}_{\mathcal{X}}}[c(\mathbf{x})=g(\mathbf{x})]\leq 1-\delta.$
@@
Definition 5.2 (Relative complexity of a function and a function class [29, Definition 6]).
@@ 5.1 Weighted Multiaccuracy and Hardcore Measures
$\min(\eta,1-\eta)+\mathrm{cor}_{\bar{\mu}}(g,\mathcal{C})\leq\max_{c\in\mathcal{C}}\Pr_{\mathbf{x}\sim\bar{\mu}}[c(\mathbf{x})=g(\mathbf{x})]\leq\max(\eta,1-\eta)+\mathrm{cor}_{\bar{\mu}}(g,\mathcal{C}).$ (10)
@@
Recall that $\bar{c}=1-c$ and that we are assuming that $\mathcal{C}$ is closed under complementation. So we have
@@
To prove the lower bound, in order to predict $g$ with accuracy $\min(\eta,1-\eta)+|\mathrm{cor}_{\bar{\mu}}(g,c)|$, we use $c$ if $\mathrm{cor}_{\bar{\mu}}(g,c)\geq 0$; else we use $\bar{c}$, which also lies in $\mathcal{C}$ by assumption. Maximizing over $c\in\mathcal{C}$ gives the claimed bound. ∎
@@
Lemma 5.7 can be seen as a generalization of Yao’s Lemma relating unpredictability with indistinguishability (see Lemma 2.14 in [8] for a formal statement).
@@
$=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{D}_{\mathcal{X}}}[c(\mathbf{x})w(p(\mathbf{x}))\cdot|g(\mathbf{x})-p(\mathbf{x})|\cdot(2g(\mathbf{x})-1)]$
@@
$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{D}_{\mathcal{X}}}[w(p(\mathbf{x}))(g(\mathbf{x})-p(\mathbf{x}))]|=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\bar{\mu}_{w}}[(2g(\mathbf{x})-1)]\cdot\mathsf{dns}(\mu)=(2\eta-1)\cdot\mathsf{dns}(\mu_{w}).$
@@
$\left|\eta-\frac{1}{2}\right|\leq\frac{\varepsilon}{2\mathsf{dns}(\mu_{w})}.$
@@
$\mathrm{cor}_{\bar{\mu}_{w}}(\mathcal{C},g)\leq\frac{\varepsilon}{\mathsf{dns}(\mu_{w})}.$
@@
$\leq\frac{1}{2}+\frac{\epsilon}{2\mathsf{dns}(\mu_{w})}+\mathrm{cor}_{\bar{\mu}_{w}}(\mathcal{C},g)$
$\leq\frac{1}{2}+\frac{3\varepsilon}{2\mathsf{dns}(\mu_{w})}$
@@ 5.2 Optimal Density Analysis Using Calibration
We will use the following identity for $v\in[0,1]$ and $y\in\{0,1\}$ which can be viewed as the multilinear expansion of |yv||y-v|:
|yv|=y(12v)+v.|y-v|=y(1-2v)+v.
Using this with $y=g(x)$ and $v=p(x)$ we can write
@@
Since $w(p)\in[1,2]$ and $1-2p\in[-1,1]$, we have $|w(p)(1-2p)|\leq 2$. Since $p$ is $\tau$-calibrated, by Equation (4) applied to v(p)=w(p)(12p)/2v(p)=w(p)(1-2p)/2, the predictor $p$ satisfies the bound
@@
[…]
@@
Often, the assumption of bounded $\mathsf{ECE}$ can be replaced with the weaker assumption that $\operatorname*{\mathbf{E}}[v(p)(y-p)]$ is bounded for specific functions $v$. In the proof of Lemma 5.10, we need to bound the calibration error for v(p)=w(p)(12p)/2v(p)=w(p)(1-2p)/2, while in the proof of Lemma 5.11, we only needed to bound the error for $v(p)=2\cdot\mathbf{1}[p\geq 1/2]-1$.
@@ 5.3 Hardcore Measures with Optimal Density
Next we consider the algorithmic complexity of obtaining a predictor that satisfies both calibration and $(w_{\mathsf{Max}},\mathcal{C})$-weighted multiaccuracy, assuming an oracle for $(\alpha,\beta)$ weak agnostic learning for $\mathcal{C}$. Following previous work, we will consider the number of invocations to the weak agnostic learner. The key point is that since the only care about a single weight function $w_{\mathsf{Max}}$ which is bounded by $2$, the complexity of the weighted versions of multiaccuracy and calibrated multiaccuracy are the same as their unweighted counterparts. The work of [17] which introduced calibrated multiaccuracy shows that its complexity is comparable to that of multiaccuracy, and much lower than multicalibration [17, Section 7]. This relies on ideas that are by now standard in the multicalibration literature [24, 19, 17], so we will provide a sketch.
@@
To achieve calibration in addition to weighted multiaccuracy, we use the analysis of the calibrated multiaccuracy algorithm from [17, Section 7]. The key observation there is that if the calibration error is large, then recalibrating the predictor reduces the squared loss. Hence we can interleave recalibration steps with the calls to the weak learner, all the while ensuring that the squared loss decreases, and the number of calls to weak learner stays bounded by $O(1/\beta^{2})$.
Lastly, we use Theorem 5.13 to derive IHCL with optimal hardcore density $2\delta$. We use the calibrated multiaccuracy algorithm from [17, Section 7] to bound the complexity $\mathcal{C}_{t,q}$ of the class on which the input function $g$ is assumed to be $\delta$-hard.
@@
Let $\mathcal{C}$ be a family of functions $c:\mathcal{X}\rightarrow\{0,1\}$, let $\mathcal{D}_{\mathcal{X}}$ be a probability distribution over $\mathcal{X}$, and let $\epsilon,\delta,\tau>0$. There exist $t=O((1/(\epsilon^{2}\delta^{2})+1/\tau)\cdot\log(|\mathcal{X}|/\epsilon))$, $q=O(1/(\epsilon^{2}\delta^{2}))$ such that the following holds: If $g:\mathcal{X}\rightarrow\{0,1\}$ is $(\mathcal{C}_{t,q},\delta)$-hard on $\mathcal{D}_{\mathcal{X}}$, then there is a measure $\mu$ satisfying:
@@
We call the $\mathsf{calMA}$ algorithm from [17] with $\mathcal{C},\mathcal{D}^{g}$, weight function $w_{\mathsf{Max}}$, desired multiaccuracy error $\epsilon\delta$, and desired calibrated error $\tau/4$, and obtain a $(w_{\mathsf{Max}},\mathcal{C},\epsilon\delta)$-multiaccurate and $(\tau/4)$-calibrated predictor $p$ for $\mathcal{D}^{g}$. Using this predictor $p$, the input function $g$, and the weight function $w_{\mathsf{Max}}$, we construct the measure $\mu_{\mathsf{Max}}$ as per Definition 5.4; namely, $\mu_{\mathsf{Max}}(x)=w_{\mathsf{Max}}(p(x))\cdot|g(x)-p(x)|$. By Theorem 5.13, this measure $\mu_{\mathsf{Max}}$ achieves the claimed hardness and optimal density stated in Theorem 5.14. For the $\mathsf{calMA}$ algorithm, we assume access to a proper $(\epsilon\delta,\epsilon\delta)$-weak agnostic learner for $\mathcal{C}$ under $\mathcal{D}_{\mathcal{X}}$; this satisfy the parameter requirements in [17, Alg. 2]. [^12] In [17], the total number of calls to an $(\alpha,\beta)$-weak agnostic learner is analyzed. Here we are not concerned with the efficiency of the learner so we can assume a strong learner giving $\alpha=\beta$.
By Theorem 7.7 in [17], the total number of calls to the learner during the execution of the $\mathsf{calMA}$ algorithm is bounded by $O(1/(\epsilon\delta)^{2})$, which yields the claimed $q$ parameter. We remark that in their algorithm, the recalibration step does not call the weak agnostic learner, which is why the calibration error $\tau/4$ is not part of the $q$ parameter. As for the total number of gates in the circuit (i.e., the $t$ parameter), each iteration of the multiaccuracy step in the $\mathsf{calMA}$ algorithm requires a scalar multiplication, a finite-precision addition, and a projection onto the interval $[0,1]$. As argued in [8, Appx. A], this can be handled by $O(\log(1/\epsilon)+\log(1/|\mathcal{X}|)$ gates, where we perform computations up to a fixed precision of $\Theta(\epsilon)$ and $\log(1/|\mathcal{X}|)$ corresponds to the bit-length of the elements in $\mathcal{X}$. Lastly, the calibration-based steps in the $\mathsf{calMA}$ algorithm do not require further calls to the learner, but they require computing all of the values of the predictor $p$ ([17, Lemma 7.4]). We discretize $p$ with parameter $\tau$, so that the interval $[0,1]$ is partitioned into $O(1/\tau)$ intervals. To compute the $O(1/\tau)$ values of the predictor $p$, we can use a lookup table of size $O(1/\tau)$; such a lookup function can be computed by a circuit of size $O(1/\tau)$ [2, §4]. ∎
@@
[8] uses the multicalibration algorithm to obtain a stronger and more general version of Impagliazzo’s hardcore lemma, IHCL++++. From it, they recover the IHCL theorem with density $2\delta$. However, because this recovery is done through the multicalibration algorithm, the $(t,q)$ parameters for the class $\mathcal{C}_{t,q}$, although polynomial in $1/\epsilon,1/\delta$, are much larger than the ones we have obtained in Theorem 5.14. In particular, using the multicalibration algorithm presented in [18], we require $q=O(1/(\epsilon^{6}\delta^{6}))$ calls to the weak agnostic learner [17]. When we recover the IHCL theorem from IHCL++++, the parameters further increase to $q=1/(\epsilon^{2}\delta)^{6}$ oracle gates. [^13] This is because we can only glue together the pieces in the partition that are large enough and balanced enough; see [8, §3.3] for details.
@@
In particular, Impagliazzo’s boosting-based proof requires $q=O(1/(\epsilon^{2}\delta^{2}))$ oracle calls [28, Thm. 1], and all subsequent hardcore constructions retain the dependence on $1/\epsilon^{2}$ in $q$ [35, 25], including the uniform versions of the hardcore lemma [47, 1, 49].
The size loss of $O(1/\epsilon^{2})$ between the class $\mathcal{C}_{t,q}$ (on which we assume the input function $g$ is mildly hard) and the class $\mathcal{C}$ (on which $g$ is strongly hard) is inevitable in all boosting-based proofs of IHCL [38], Blanc, Hayderi, Koch, and Tan recently showed that this circuit size loss is necessary and that these parameters are in fact optimal and inherent to the hardcore lemma, regardless of proof strategy [4]. The dependence on $\delta$ can be improved: Nisan’s min-max proof in Impagliazzo’s original paper yields $q=O(1/\epsilon^{2}\log(1/(\epsilon\delta))$ [28, Lemma 2], and it was further improved to $q=O(1/\epsilon^{2}\log(1/\delta))$ by Kilvans and Servedio who showed an elegant connection to boosting [35, Thm. 25]. But these proofs do not give the optimal density.
Deriving IHCL from IHCL++++.
The IHCL++++ theorem of [8] partitions the space $\mathcal{X}$ into level sets $S_{v}$. Within each level set, they construct a hardcore $\mathcal{H}_{v}$ distribution of density $2\min(v,1-v)$. We also know that $\operatorname*{\mathbf{E}}[\min(v,1-v)]\geq\delta$. To derive the IHCL with density $2\delta$, we need to glue these distributions together to get a measure of density $2\delta$. In their paper, [8] suggest the following gluing: sample $S_{v}$ according to $\mathcal{D}_{\mathcal{X}}$ and then sample within $S_{v}$ from $\mathcal{H}_{v}$. This gluing as stated might not give a distribution with bounded density. The reason being that when we take convex combinations of distributions over disjoint regions, the min-entropy of the combination is the maximum of the min-entropies, not the average. A simple fix is to go from hardcore distributions to sets, and then take the union of sets.